OpenStudy (anonymous):

the difference between six times a number and 9 is equal to five times the sum of the number and 2. Find the number.

OpenStudy (anonymous):

"six times a number" = 6x "the difference between six times a number and 9" = 6x-9 "sum of the number and 2" = x + 2 "five times the sum of the number and 2" = 5(x+2).

OpenStudy (anonymous):

6x-9=5(x+2)

OpenStudy (anonymous):

Bingo! Solve for x.

OpenStudy (anonymous):

how do i solve that?

OpenStudy (anonymous):

x=19!

OpenStudy (anonymous):

can you show me how you did that plz?

OpenStudy (anonymous):

It's basic algebra from there. First step is to use the distributive property: \[a(b+c)=ab+ac\]\[\Rightarrow 5(x+2)=5x+5(2)\]

OpenStudy (anonymous):

So \[6x-9=5x+10\]Take 5x from both sides:\[6x-5x-9=5x-5x+10\]\[\Leftrightarrow x-9=10\]Add 9 to both sides:\[x-9+9 = 10+9\]\[\Leftrightarrow x=19\]

OpenStudy (anonymous):

right what Dalvoron said ^.